# FORTALESA: Fault-Tolerant Reconfigurable Systolic Array for DNN Inference

### v1

@@ Abstract
The emergence of Deep Neural Networks (DNNs) in mission- and safety-critical applications brings their reliability to the front. High performance demands of DNNs require the use of specialized hardware accelerators. Systolic array architecture is widely used in DNN accelerators due to its parallelism and regular structure. This work presents a run-time reconfigurable systolic array architecture with three execution modes and four implementation options. All four implementations are evaluated in terms of resource utilization, throughput, and fault tolerance improvement. The proposed architecture is used for reliability enhancement of DNN inference on systolic array through heterogeneous mapping of different network layers to different execution modes. The approach is supported by a novel reliability assessment method based on fault propagation analysis. It is used for the exploration of the appropriate execution mode–layer mapping for DNN inference. The proposed architecture efficiently protects registers and MAC units of systolic array PEs from transient and permanent faults. The reconfigurability feature enables a speedup of up to 3×3$\times$3 ×, depending on layer vulnerability. Furthermore, it requires 6×6$\times$6 × less resources compared to static redundancy and 2.5×2.5$\times$2.5 × less resources compared to the previously proposed solution for transient faults.
Index Terms:
systolic array, fault tolerance, deep neural networks, accelerator.
@@ I Introduction
Recent years have been marked by a tremendous increase in adopting Machine Learning (ML) algorithms, specifically Deep Neural Networks (DNNs), and their employment in mission- and safety-critical application domains, such as automotive, avionics, and space. The computational complexity of DNN models requires the use of specialized hardware accelerators. Many DNN accelerators are based on systolic array architecture, including Google TPU [1], MIT Eyeriss [2], and Berkeley Gemmini [3]. However, as nanoelectronics used in hardware accelerators are prone to faults caused by radiation and aging, the concern about their reliability gets to the front [4].
Furthermore, with the current need for self-aware computing systems that can monitor their state and adapt to changing environments and operating conditions, an important property of hardware design is the ability to adjust or reconfigure to changing conditions for improved availability. To address those needs, we propose a run-time reconfigurable systolic array architecture that tackles the balance between reliability and performance through different execution modes. As a target application, we consider the inference of DNNs. Since different DNN components have different vulnerability levels, they might benefit from heterogeneous mapping to different execution modes of the proposed systolic array. We also introduce a reliability assessment method to efficiently validate the effectiveness of the proposed architecture and determine the appropriate execution mode–layer mapping.
@@
  • a novel reliability assessment method for DNN inference on systolic array based on fault propagation analysis;
@@
The proposed architecture efficiently protects both sequential (registers) and combinational (MAC) components of systolic array processing elements. While the main objective is to address transient faults (soft errors), the proposed architecture can handle permanent faults as well allowing for graceful degradation. The reconfigurability feature of FORTALESA enables a speedup up to 3×3$\times$3 ×, depending on layer vulnerability, and requires 6×6$\times$6 × less resources compared to static redundancy and 2.5×2.5$\times$2.5 × less resources than the previously proposed solution for transient faults.
The rest of the paper is organized the following way. Section II gives an overview of related works. Section III presents the background on systolic array architecture. Section IV introduces the architecture of the proposed reconfigurable systolic array. Section V explains the proposed reliability assessment method based on fault propagation. Section VI evaluates the effectiveness of the proposed methodology for DNN reliability, and Section VII concludes the paper.
@@ II Related works
Systolic arrays were first introduced in the 1980s [21] and recently became popular for accelerating DNNs. As the reliability of DNN accelerators comes to the front, recent works have proposed methods for the assessment and enhancement of systolic arrays. [22] analyzed the effects of transient faults in the combinational logic of systolic arrays. [23] presented Saca-FI, a microarchitecture-level fault injection framework for systolic array-based CNN accelerators, for the analysis of transient faults in registers. Saca-FI is based on SCALE-Sim, a cycle-accurate simulator for DNN inference on systolic arrays [24]. [25] proposed an RTL-level fault injection framework for the analysis of stuck-at faults within multiply and accumulate (MAC) units. [26] presented a fault injection framework that models systolic array dataflow using Uniform Recurrent Equations. Unlike previous works, the proposed reliability assessment method is based on fault propagation analysis. Instead of injecting faults in the microarchitectural or RTL model of a systolic array, which is a very time-consuming task; errors are added directly to the DNN layer output using fault propagation analysis.
Reliability enhancement methods for systolic arrays, similarly to DNNs, include ABFT and network retraining. [27] presented a permanent fault correction technique. [28] proposed fault-aware pruning plus retraining (FAP+T) technique to address manufacturing defects in the systolic array-based accelerators. It requires modification of the target DNN architecture through pruning and retraining to adapt it to a specific defect in the chip. [29] presented an ABFT technique for permanent fault detection in systolic arrays on FPGAs. Permanent faults in configuration memory are considered in that work. Another permanent fault detection approach assuming faults in registers and MAC units is presented in [30]. The approach is based on checksum calculations and was implemented on FPGA as well. While most works address permanent faults, [31] proposed an ABFT method for detecting timing faults due to reduced voltage. Compared to the aforementioned works, the proposed reconfigurable systolic array architecture addresses both permanent and transient faults and does not require retraining of a DNN for each particular chip.
@@ III-A Systolic array architecture
Systolic array has a regular structure consisting of an array of Processing Elements (PEs). Each PE consists of a Multiply-Accumulate (MAC) unit and registers to hold the intermediate values. Memory buffers for activations and weights are located at the top and left sides of the systolic array correspondingly. Commonly, weights are fed vertically from top to bottom, thus each column is assigned to one output channel, and activations (inputs) are fed from left to right, hence each row processes values from the same sliding windows. Dataflow type defines the internal structure of PEs and the movement of data through the array. Dataflow types include weight-stationary (WS), input-stationary (IS), and output-stationary (OS) [36].
@@
In the input-stationary dataflow, activations are pre-loaded in the systolic array in a similar manner as WS and kept in the systolic array during the matrix multiplication process. The data movement and structure of the systolic array are similar to WS dataflow.
In the output-stationary dataflow, outputs are kept in the PEs of the systolic array, while weights and activations are streaming through the systolic array (Fig. 1). Weights are streaming from top to bottom, and activations (inputs) are streaming from left to right. Outputs are accumulated inside PEs (Fig. 2). This dataflow type does not need a pre-loading stage. An output-stationary systolic array is considered for this work. It is commonly used and allows high inputs and outputs reuse [37, 38] and on average has higher performance [33]. Previous works have shown that it is more fault tolerant [39].
Figure 1: Dataflow in the output-stationary systolic array
Figure 2: Architecture of the individual PE in the output-stationary systolic array
@@
where $T_{a}$ and $T_{w}$ are the number of tiles in activations and weights correspondingly, $P$ is the size of the reshaped activations, and $K$ is the number of output channels and the size of the weights matrix. The total amount of steps needed to calculate the output equals $S=T_{a}\cdot T_{w}$, and the total latency of matrix multiplication on the systolic array is therefore
@@
III-B Fault propagation in systolic arrays
Considering the structure of a single processing element, faults can occur in registers holding intermediate values or in a multiplier. Considering the regular structure of a systolic array and data movement, faults in different components produce different error patterns in the output. Fig. 3 presents the possible patterns for the output-stationary systolic array. A bit flip in a weight register (WREG) results in the propagation of faulty weight through the whole column affecting several output values in one channel. This ends up in the line pattern. A bit flip in an input register (IREG) results in the propagation of faulty activation through the whole row affecting one value in several output channels. This ends up in the bullet pattern. A bit flip in the multiplier or an output register keeping partial sum (OREG) affects one value in one output channel resulting in a point pattern.
@@ IV Fault-tolerant reconfigurable systolic array
We propose FORTALESA, a systolic array with reconfigurable redundancy that provides flexible reliability performance trade-off for critical tasks requiring fault tolerance and non-critical ones that benefit from increased parallelism.
@@
  • dual modular redundancy for reliability performance trade-off;
  • triple modular redundancy for full protection.
The selected mode is configured through a control signal that controls multiplexers. The control signal is assumed to come from the host processor based on the task requirement. Since different layers of DNNs have different vulnerability, it is possible to execute different layers in different modes, thus achieving an efficient trade-off between performance and reliability.
DMR execution mode. In the DMR mode, two neighboring PEs form a group (Fig. 4). They receive the same inputs, one of them acts as a main PE, the other as a shadow. Main PE compares two partial sums (OREG) and adjusts its partial sum value without re-execution using one of possible fault correction techniques: (a) averaging of two values (DMRA) or (b) setting faulty bits to zero (DMR0). The choice of the fault correction technique is a design-time parameter that defines two implementation options for the DMR mode.
Figure 4: Arrangement of the systolic array in the dual modular redundancy execution mode (DMRA or DMR0)
Additional connections are introduced to support redundant modes. To ensure that faulty IREG or WREG values do not propagate through the array, each PE passes the activation to the next PE of the same type skipping one in between. The shadow PE also passes the value of the OREG (partial sum) to the main PE for comparison. Main PE has additional functionality for fault correction that works in parallel with the MAC unit. Since the updated partial sum is available on the next clock cycle, the latency of the systolic array working in DMR mode equals the latency of the $N\times N/2$ systolic array plus 1:
$L_{SA}^{dmr}=M+N-1+\frac{N}{2}-1+1=M+\frac{3N}{2}-1\,.$ (5)
@@
$L_{total}^{dmr}=\left\lceil\frac{P}{N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot\left(M+\frac{3N}{2}-1\right).$ (6)
Since in the redundancy execution modes, several PEs calculate the same values, the term effective size of the systolic array is introduced. It defines the number of groups calculating unique values and the size of the output matrix. The effective size of the systolic array in the DMR execution mode is $N\times\frac{N}{2}$.
TMR execution mode. For the TMR execution mode, two implementations are possible: with groups of three or four PEs. The implementation option with groups of three PEs (TMR3) is shown in Fig. 5 and the implementation option with groups of four PEs (TMR4) in Fig. 6. The group size is the design-time parameter for the TMR mode.
Figure 5: Arrangement of the systolic array in the triple modular redundancy execution mode, implementation option with groups of three PEs (TMR3)
Figure 6: Arrangement of the systolic array in the triple modular redundancy execution mode, implementation option with groups of four PEs (TMR4)
In the TMR3 implementation, the voter in the main PE works in parallel with the MAC unit. In the TMR4 implementation, the MAC unit in the main PE is not used in the TMR mode, the main PE only compares the partial sums calculated by three shadow PEs.
The effective size of the systolic array, i.e., the size of the output matrix, for TMR3 implementation is $\frac{2N}{3}\times\frac{N}{2}$. The effective sizes of the matrices that can be multiplied are $\frac{2N}{3}\times M$ and $M\times\frac{N}{2}$. The latency of the TMR3 implementation is
$L_{SA}^{tmr3}=M+\frac{2N}{3}-1+\frac{N}{2}-1+1=M+\frac{7N}{6}-1\,.$ (7)
@@
$L_{total}^{tmr3}=\left\lceil\frac{3P}{2N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot\left(M+\frac{7N}{6}-1\right).$ (8)
The effective size of the systolic array for the TMR4 implementation is $\frac{N}{2}\times\frac{N}{2}$. The effective sizes of the matrices that can be multiplied are $\frac{N}{2}\times M$ and $M\times\frac{N}{2}$. The latency of the TMR4 implementation is, therefore,
$L_{SA}^{tmr4}=M+\frac{N}{2}-1+\frac{N}{2}-1+1=M+N-1\,.$ (9)
@@
$L_{total}^{tmr4}=\left\lceil\frac{2P}{N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot(M+N-1).$ (10)
Since both execution modes have design-time parameters, four implementation options are possible: PM-DMR0-TMR3, PM-DMR0-TMR4, PM-DMRA-TMR3, and PM-DMRA-TMR4. Their efficiency is evaluated in Section VI. The overview of the execution modes and implementation options is given in Table I.
TABLE I: Execution modes and implementation options
@@
DMR DMRA $\frac{N}{2}\times N$
DMR0
TMR TMR3 $\frac{2N}{3}\times\frac{N}{2}$
TMR4 $\frac{N}{2}\times\frac{N}{2}$
V Reliability assessment
This section presents the proposed reliability assessment method. First, for the analysis of transient and permanent faults in the regular OS systolic array used for acceleration of DNNs. Then, for assessment of the proposed architecture since it introduces the DMR execution mode with fault correction. Lastly, it explains the implementation of the assessment method.
The proposed reliability assessment method combines fault injection with fault propagation analysis. Instead of modeling a systolic array on the microarchitecture level, fault propagation analysis is used to calculate the resulting error in the layer output.
V-A Fault propagation analysis for transient faults

### v2

@@ Abstract
The emergence of Deep Neural Networks (DNNs) in mission- and safety-critical applications brings their reliability to the front. High performance demands of DNNs require the use of specialized hardware accelerators. Systolic array architecture is widely used in DNN accelerators due to its parallelism and regular structure. This work presents a run-time reconfigurable systolic array architecture with three execution modes and four implementation options. All four implementations are evaluated in terms of resource utilization, throughput, and fault tolerance improvement. The proposed architecture is used for reliability enhancement of DNN inference on systolic array through heterogeneous mapping of different network layers to different execution modes. The approach is supported by a novel reliability assessment method based on fault propagation analysis. It is used for the exploration of the appropriate execution mode–layer mapping for DNN inference. The proposed architecture efficiently protects registers and MAC units of systolic array PEs from transient and permanent faults. The reconfigurability feature enables a speedup of up to $3\times$, depending on layer vulnerability. Furthermore, it requires $6\times$ fewer resources compared to static redundancy and $2.5\times$ fewer resources compared to the previously proposed solution for transient faults.
@@ I Introduction
Recent years have been marked by a tremendous increase in adoption of Machine Learning (ML) algorithms, specifically Deep Neural Networks (DNNs), and their employment in mission- and safety-critical application domains, such as automotive, avionics, and space. The computational complexity of DNN models requires the use of specialized hardware accelerators. Many DNN accelerators are based on systolic array architecture, including Google TPU [1], MIT Eyeriss [2], and Berkeley Gemmini [3]. However, as nanoelectronics used in hardware accelerators are prone to faults caused by radiation and aging, the concern about their reliability comes to the front [4].
Furthermore, with the current need for self-aware computing systems that can monitor their state and adapt to changing environments and operating conditions, an important property of hardware design is the ability to adjust or reconfigure to changing conditions for improved availability. To address those needs, we propose a run-time reconfigurable systolic array architecture that tackles the balance between reliability and performance through different execution modes. As a target application, we consider the inference of DNNs. Since different DNN components have different vulnerability levels, they might benefit from heterogeneous mapping to different execution modes of the proposed systolic array. We also introduce a novel reliability assessment method for systolic arrays. It is used to determine the appropriate execution mode–layer mapping and consequently to validate the effectiveness of the proposed architecture.
@@
  • a novel reliability assessment method for DNN inference on a systolic array based on fault propagation analysis;
@@
The proposed architecture efficiently protects both sequential (registers) and combinational components of systolic array processing elements. While the main objective is to address transient faults (soft errors), the proposed architecture can handle permanent faults as well, allowing for graceful degradation. The reconfigurability feature of FORTALESA enables a speedup of up to $3\times$, depending on layer vulnerability, and requires $6\times$ fewer resources compared to static redundancy and $2.5\times$ fewer resources than the previously proposed solution for transient faults.
The rest of the paper is organized the following way. Section II gives an overview of related works. Section III presents the background on systolic array architecture. Section IV introduces the architecture of the proposed reconfigurable systolic array. Section V explains the proposed reliability assessment method for systolic arrays based on fault propagation. Section VI evaluates the effectiveness of the proposed methodology for DNN reliability, and Section VII concludes the paper.
@@ II Related works
Systolic arrays were first introduced in the 1980s [21] and recently became popular for accelerating DNNs. As the reliability of DNN accelerators comes to the front, recent works have proposed methods for the assessment and enhancement of systolic arrays. [22] analyzed the effects of transient faults in the combinational logic of systolic arrays. [23] presented Saca-FI, a microarchitecture-level fault injection framework for systolic array-based CNN accelerators, for the analysis of transient faults in registers. Saca-FI is based on SCALE-Sim, a cycle-accurate simulator for DNN inference on systolic arrays [24]. [25] proposed an RTL-level fault injection framework for the analysis of stuck-at faults within multiply and accumulate (MAC) units. [26] presented a fault injection framework that models systolic array dataflow using Uniform Recurrent Equations. Unlike previous works, the proposed reliability assessment method is based on fault propagation analysis. Instead of injecting faults in the microarchitectural or RTL model of a systolic array, which is a very time-consuming task; errors are added directly to the DNN layer output using fault propagation analysis that allows to significantly speed up the fault injection process.
Reliability enhancement methods for systolic arrays, similarly to DNNs, include ABFT and network retraining. [27] presented a permanent fault correction technique by introducing a re-computing unit that recalculates the computations assigned to faulty processing elements. [28] addressed permanent faults through bypassing faulty units and assigning their computations to the redundant ones. Similarly, [29] used pruning of faulty processing elements for permanent fault correction. [30] proposed a fault-aware pruning plus retraining (FAP+T) technique to address manufacturing defects in the systolic array-based accelerators. It requires modification of the target DNN architecture through pruning and retraining to adapt it to a specific defect in the chip. [31] presented an ABFT technique for permanent fault detection in systolic arrays on FPGAs. Permanent faults in configuration memory are considered in that work. Another permanent fault detection approach assuming faults in registers and MAC units is presented in [32]. The approach is based on checksum calculations and was implemented on FPGA as well. [33] presented a fault tolerance solution for an FPGA-based systolic array accelerator. Faults in configuration memory affecting the systolic array datapath are detected using checksum values and corrected using dynamic partial reconfiguration of the FPGA. [34, 35] presented an ABFT methodology specifically developed for sparse systolic arrays, considering their unique characteristics. While most works address permanent faults, [36] proposed an ABFT method for detecting timing faults due to reduced voltage. Several works proposed test architectures for permanent and transition-delay fault detection in low-power accelerators based on systolic arrays [37, 38]. Compared to the aforementioned works that mostly address detection of permanent faults, the proposed reconfigurable systolic array architecture addresses both permanent and transient faults, enables reliability-performance trade-off, provides fault correction in both fault-tolerant execution modes, and does not require retraining of a DNN for each particular chip.
@@ III-A Systolic array architecture
Systolic array has a regular structure consisting of an array of Processing Elements (PEs). Each PE consists of a Multiply-Accumulate (MAC) unit and registers to hold the intermediate values. Memory buffers for activations and weights are located at the top and left sides of the systolic array respectively. Commonly, weights are fed vertically from top to bottom, thus each column is assigned to one output channel, and activations (inputs) are fed from left to right, hence each row processes values from the same sliding windows. Dataflow type defines the internal structure of PEs and the movement of data through the array. Dataflow types include weight-stationary (WS), input-stationary (IS), and output-stationary (OS) [43].
@@
In the input-stationary dataflow, activations are pre-loaded in the systolic array similarly to WS and kept in the systolic array during the matrix multiplication process. The data movement and structure of the systolic array are similar to WS dataflow.
In the output-stationary dataflow, outputs are kept in the PEs of the systolic array, while weights and activations are streaming through the systolic array. Weights are streaming from top to bottom, and activations (inputs) are streaming from left to right. Outputs are accumulated inside PEs. This dataflow type does not need a pre-loading stage.
An output-stationary systolic array is considered for this work. It is commonly used and allows high inputs and outputs reuse [44, 45] and, on average, has higher performance [40]. Previous works have shown that it is more fault-tolerant [46].
III-B Convolution on systolic array
@@
where $T_{a}$ and $T_{w}$ are the number of tiles in activations and weights correspondingly, $P$ is the size of the reshaped activations, and $K$ is the number of output channels and the size of the weights matrix. The total number of steps needed to calculate the output equals $S=T_{a}\cdot T_{w}$, and the total latency of matrix multiplication on the systolic array is therefore
@@ IV Fault-tolerant reconfigurable systolic array
We propose FORTALESA, a systolic array with reconfigurable redundancy that provides flexible reliability–performance trade-off for critical tasks requiring fault tolerance and non-critical ones that benefit from increased parallelism.
@@
  • dual-redundancy grouping (DRG) for reliability–performance trade-off;
  • triple-redundancy grouping (TRG) for full protection.
The selected execution mode is configured through a control signal that controls multiplexers. The control signal comes from the host processor based on the task requirement. Since different layers of DNNs have different vulnerability, it is possible to execute different layers in different modes, thus achieving an efficient trade-off between performance and reliability.
The proposed architecture addresses faults in registers and arithmetic units of systolic array PEs, since they lead to silent data corruption without indication of an error, unlike faults in control logic that usually result in functional interrupts that signal that there was a fault and require a reset.
DRG execution mode. In the DRG execution mode, two neighboring PEs form a group (Fig. 1). They receive the same inputs, one of them acts as the main PE, the other as a shadow. Main PE compares two partial sums (OREG) and adjusts its partial sum value without re-execution using one of the possible fault correction techniques: (a) averaging of two values (DRGA) or (b) setting faulty bits to zero (DRG0). While dual redundancy does not enable full protection, these correction techniques decrease the effect of the fault on the output values, thus enabling a reliability–performance trade-off. The choice of the fault correction technique is a design-time parameter that defines two implementation options for the DRG execution mode.
Figure 1: Arrangement of the systolic array in the dual-redundancy grouping execution mode (DRGA or DRG0)
Additional connections are introduced to support redundant execution modes. To ensure that faulty IREG or WREG values do not propagate through the array, each PE passes the activation to the next PE of the same type, skipping one in between. While faults in buffers and interconnects are not in the focus of this work, such interleaved connections can help protect against those faults as well. During the fault-tolerant execution modes, there are two copies of input data traversing through the systolic array. The shadow PE also passes the value of the OREG (partial sum) to the main PE for comparison. Main PE has additional functionality for fault correction that works in parallel with the MAC unit. Since the updated partial sum is available on the next clock cycle, the latency of the systolic array working in DRG execution mode equals the latency of the $N\times N/2$ systolic array plus 1:
$L_{SA}^{drg}=M+N-1+\frac{N}{2}-1+1=M+\frac{3N}{2}-1\,.$ (5)
@@
$L_{total}^{drg}=\left\lceil\frac{P}{N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot\left(M+\frac{3N}{2}-1\right).$ (6)
Since in the redundancy execution modes, several PEs calculate the same values, the term effective size of the systolic array is introduced. It defines the number of groups calculating unique values and the size of the output matrix. The effective size of the systolic array in the DRG execution mode is $N\times\frac{N}{2}$.
TRG execution mode. For the TRG execution mode, two implementations are possible: with groups of three or four PEs. The implementation option with groups of three PEs (TRG3) is shown in Fig. 2 and the implementation option with groups of four PEs (TRG4) in Fig. 3. The group size is the design-time parameter.
Figure 2: Arrangement of the systolic array in the triple-redundancy grouping execution mode, implementation option with groups of three PEs (TRG3)
Figure 3: Arrangement of the systolic array in the triple-redundancy grouping execution mode, implementation option with groups of four PEs (TRG4)
In the TRG3 implementation, the voter in the main PE works in parallel with the MAC unit. In the TRG4 implementation, the MAC unit in the main PE is not used in the TRG execution mode, the main PE only compares the partial sums calculated by three shadow PEs.
The effective size of the systolic array, i.e., the size of the output matrix, for TRG3 implementation is $\frac{2N}{3}\times\frac{N}{2}$. The effective sizes of the matrices that can be multiplied are $\frac{2N}{3}\times M$ and $M\times\frac{N}{2}$. The latency of the TRG3 implementation is
$L_{SA}^{trg3}=M+\frac{2N}{3}-1+\frac{N}{2}-1+1=M+\frac{7N}{6}-1\,.$ (7)
@@
$L_{total}^{trg3}=\left\lceil\frac{3P}{2N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot\left(M+\frac{7N}{6}-1\right).$ (8)
The effective size of the systolic array for the TRG4 implementation is $\frac{N}{2}\times\frac{N}{2}$. The effective sizes of the matrices that can be multiplied are $\frac{N}{2}\times M$ and $M\times\frac{N}{2}$. The latency of the TRG4 implementation is, therefore,
$L_{SA}^{trg4}=M+\frac{N}{2}-1+\frac{N}{2}-1+1=M+N-1\,.$ (9)
@@
$L_{total}^{trg4}=\left\lceil\frac{2P}{N}\right\rceil\cdot\left\lceil\frac{2K}{N}\right\rceil\cdot(M+N-1).$ (10)
Since both execution modes have design-time parameters, four implementation options are possible: PM-DRG0-TRG3, PM-DRG0-TRG4, PM-DRGA-TRG3, and PM-DRGA-TRG4. Their efficiency is evaluated in Section VI. The overview of the execution modes and implementation options is given in Table I.
TABLE I: Execution modes (run-time) and implementation options (design-time)
@@
DRG DRGA $\frac{N}{2}\times N$
DRG0
TRG TRG3 $\frac{2N}{3}\times\frac{N}{2}$
TRG4 $\frac{N}{2}\times\frac{N}{2}$
The architecture of FORTALESA is based on the output-stationary dataflow; however, other dataflows can be used to build reconfigurable systolic arrays following FORTALESA methodology. Processing elements can be grouped in a similar way, with similar interleaved connections. During the pre-loading stage, PEs from one group should receive the same input or weight. The faults would be corrected by applying a selected correction technique to the output value.
V Reliability assessment method for systolic arrays
This section presents the proposed reliability assessment method for systolic arrays. The proposed reliability assessment method combines fault injection with fault propagation analysis. Instead of modeling a systolic array on the microarchitecture level and performing time-consuming cycle-accurate simulations, fault propagation analysis is used to calculate the resulting error in the layer output. This allows to noticeable speed up the reliability assessment of DNN inference on a systolic array. The method allows to analyze both transient and permanent faults in the OS systolic arrays.
V-A Fault propagation in systolic arrays
Considering the structure of a single processing element, faults can occur in registers holding intermediate values or in a multiplier. Considering the regular structure of a systolic array and data movement, faults in different components produce different error patterns in the output tensor. Fig. 4 presents the possible patterns for the convolution operation on the output-stationary systolic array. Fault patterns are derived analytically and experimentally following examples from previous works for similar cases [31, 48]. A bit flip in a weight register (WREG) results in the propagation of faulty weight through the whole column affecting several output values in one channel. This ends up in the line pattern. A bit flip in an input register (IREG) results in the propagation of faulty activation through the whole row affecting one value in several output channels. This ends up in the bullet pattern. A bit flip in the multiplier or an output register keeping partial sum (OREG) affects one value in one output channel, resulting in a point pattern.
V-B Fault propagation analysis for transient faults